OpenStudy (anonymous):

Please, just one person at a time. log(2)256 The base is two. How do you solve this? This is what I did: log(2)256 = x 2^x = 256 x = 8 Did I do this right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

yes

OpenStudy (campbell_st):

rewrite 256 as a power of 2 well 256 = 2^8 so you are looking as solving \[\log_{2}(2^8) = x\] so applying some log laws \[\log(x^a) = a \times \log(x)\] and when the base of the log and exponent are the same \[\log_{a}(a) = 1\] so now apply these laws to your question in the rewritten form to find the value of x

OpenStudy (anonymous):

Thank you